# Shower formation in the presence of a string-inspired foam in space-time

### v1

@@ 3 Instead of conclusions
Our modest proposal [19, 20, 21] that a possible signature of light-speed variation in cosmic photons could be attributed to propagation in a stochastic QG D-foam medium, stimulated us to examine effects induced in atmospheric showers of these photons. The existence of “shower anomalies” has been previously suggested based on general form of Planck scale-motivated LV dispersion relations which characterize the effective field theory model of LV in [37, 36]. Using the concept of radiation formation length we have shown, however, that such anomalies may be eliminated in the present model. Qualitatively, this difference has been traced to the fact that the assumption of energy conservation implicit in previous analyses of scattering kinematics does not hold in foamy situations. Our lack of knowledge of D-foam/matter interaction dynamics for which presently one does not have analytic control – a major limitation for conclusions drawn from above – is of much smaller impact as we did tentative calculations while such aspects nevertheless deserve future investigation.
@@
In any rate, as a cautionary theoretical note, one always needs a concrete formalism to study possible effects for the relativistic kinematics. This is in contrast to the result [32] that both the deformed velocity of light and the suppressed photon showers can all be induced by the same subluminal photon dispersion (1). This underlines the limitations of a purely phenomenological model: without explicit guidance from a QG theory it is difficult to be certain of the possible LV in different observational contexts.
@@
Note that while another process, the direct photonuclear reaction, can also be $\gamma$-rays’ first reaction in air, it is much inefficient than BH ones even for photon at $10^{19}$ eV, which has not been detected yet. Still, we want to give some clues to the problem of whether it and neutrino–nucleus interaction are affected, since at least the latter is relevant for detection of LV cosmic neutrinos (3[47, 48, 49], at IceCube [50], a Cherenkov detector. We note, again, that it is a prerequisite of foam corrections that there is nontrivial interaction between the incident particles with D-foam. The fact that atomic nucleus consists of protons and neutrons which are not structureless particles complicates matters. Although the neutrons are supposed to interact with this foam, their quark substructure may suppress this by introducing additional (and significant) suppression factors in the pertinent QG parameter, due to electric charge of the latter, thereby leading to hardly modified neutrino (or photon) scattering off a nucleus.
@@
We should emphasize, however, that at the present level of understanding of the model(s) describing the recoil foam backgrounds one is unable to rigorously establish that neutrino initiated cascade within the detector material is unaffected (or, negligibly affected) by the new features. But, as discussed, the nonelementary quark and gluon structure of nuclei does play an important suppressing role of any foam effects. Since the D-foam scenario addressed here has been found to introduce rather smooth departures from relativity (smoother than in a LV theory), it is natural to expect, that nothing much would change for such kind of phenomena involving particles of a few $10^{15}$ eV and even the recent 220 PeV KM3NeT neutrino [51] (still small in comparison to the scale characteristic of any QG-motivated effects for a detector). In fact Lorentz symmetry is restored on average in this type of models by the zero mean of the distribution of $r$ despite the presence of recoil $u\sim g_{s}rE/M_{s}$, as pointed out elsewhere [21] for other purposes.
@@
A final remark is that nonconservation of particle energy that is due to background changes during particle interactions in D-brane foam does not make its appearance during propagation, where momentum of an observable particle is exactly conserved [52, 53], following the closure of a worldsheet logarithmic algebra, and, its energy conserved on the average [23] (see also first review in [1]). For photon propagation in the vacuum of D-foam space-time, one would not expect any uncertainties in energies and momenta to dominate over the modifications of the dispersion relations (4). This is to be contrasted with foam cases like that of [54, 55] where energy–momentum fluctuations dominate the propagation and are bounded by analyzing cosmic-ray data. In our approach, energy violation is merely relevant for interactions of $\gamma$-rays with atmospheric nuclei as discussed here or with cosmic background fields. The foam impact on the latter reaction is up for debate, and, as already mentioned, is left for future study.
The main conclusion from this Letter, therefore, is that, in the string model of quantum foam considered here, subluminal modified photon dispersion relations (4) that may be compatible with a phenomenological hint of light-speed variation at a few $10^{17}$ GeV [7, 8, 9, 5, 6, 10, 11, 12, 13, 14, 16, 17, 18, 15] can evade drastic modifications to pair creation and bremsstrahlung under certain conditions. Thus, LV with such intensities may concordant in the context of such models with tests of subluminal photons in giant air showers. Otherwise the strong suppression effects induced in other LV proposals that are not observed experimentally in shower formation would allow [36] to imply trans-Planckian limits for the relevant scale. A detailed estimate of more complicated effects of foam from dynamical aspect is anticipated to refine the above analysis, which might help better probing or constraining this type of QG models from string theory.
@@ Acknowledgments
A potential new line of evidence for the light-speed variation [7, 8, 9, 5, 6, 10, 11, 12, 13, 14, 16, 17, 18, 15] was recently reported [56], based on the detection of a 300 TeV photon-like air shower event by Carpet–3 coincident with GRB 221009A, offering further support for the theory here championed.

### v2

@@ 3 Instead of conclusions
Our modest proposal Li:2021gah , Li:2021tcw , Li:2022sgs that a possible signature of light-speed variation in cosmic photons could be attributed to propagation in a stochastic QG D-foam medium, stimulated us to examine effects induced in atmospheric showers of these photons. The existence of “shower anomalies” has been previously suggested based on general form of Planck scale-motivated LV dispersion relations which characterize the effective field theory model of LV in Myers:2003fd , Satunin:2023yvj . Using the concept of radiation formation length we have shown, however, that such anomalies may be eliminated in the present model. Qualitatively, this difference has been traced to the fact that the assumption of energy conservation implicit in previous analyses of scattering kinematics does not hold in foamy situations. Our lack of knowledge of D-foam/matter interaction dynamics for which presently one does not have analytic control – a major limitation for conclusions drawn from above – is of much smaller impact as we did tentative calculations while such aspects nevertheless deserve future investigation.
@@
In any rate, as a cautionary theoretical note, one always needs a concrete formalism to study possible effects for the relativistic kinematics. This is in contrast to the result Vankov:2002gt that both the deformed velocity of light and the suppressed photon showers can all be induced by the same subluminal photon dispersion (1). This underlines the limitations of a purely phenomenological model: without explicit guidance from a QG theory it is difficult to be certain of the possible LV in different observational contexts.
@@
Note that while another process, the direct photonuclear reaction, can also be $\gamma$-rays’ first reaction in air, it is much inefficient than BH ones even for photon at $10^{19}$ eV, which has not been detected yet. Still, we want to give some clues to the problem of whether it and neutrino–nucleus interaction are affected, since at least the latter is relevant for detection of LV cosmic neutrinos (3Huang:2018ham , Huang:2019etr , Huang:2022xto , at IceCube IceCube:2013cdg , a Cherenkov detector. We note, again, that it is a prerequisite of foam corrections that there is nontrivial interaction between the incident particles with D-foam. The fact that atomic nucleus consists of protons and neutrons which are not structureless particles complicates matters. Although the neutrons are supposed to interact with this foam, their quark substructure may suppress this by introducing additional (and significant) suppression factors in the pertinent QG parameter, due to electric charge of the latter, thereby leading to hardly modified neutrino (or photon) scattering off a nucleus.
@@
We should emphasize, however, that at the present level of understanding of the model(s) describing the recoil foam backgrounds one is unable to rigorously establish that neutrino initiated cascade within the detector material is unaffected (or, negligibly affected) by the new features. But, as discussed, the nonelementary quark and gluon structure of nuclei does play an important suppressing role of any foam effects. Since the D-foam scenario addressed here has been found to introduce rather smooth departures from relativity (smoother than in a LV theory), it is natural to expect, that nothing much would change for such kind of phenomena involving particles of a few $10^{15}$ eV and even the recent 220 PeV KM3NeT neutrino KM3NeT:2025npi  (still small in comparison to the scale characteristic of any QG-motivated effects for a detector). In fact Lorentz symmetry is restored on average in this type of models by the zero mean of the distribution of $r$ despite the presence of recoil $u\sim g_{s}rE/M_{s}$, as pointed out elsewhere Li:2022sgs for other purposes.
@@
A final remark is that nonconservation of particle energy that is due to background changes during particle interactions in D-brane foam does not make its appearance during propagation, where momentum of an observable particle is exactly conserved Kogan:1996zv , Mavromatos:1998nz , following the closure of a worldsheet logarithmic algebra, and, its energy conserved on the average Ellis:2000sf  (see also first review in Li:2025stf ). For photon propagation in the vacuum of D-foam space-time, one would not expect any uncertainties in energies and momenta to dominate over the modifications of the dispersion relations (4). This is to be contrasted with foam cases like that of Amelino-Camelia:2002qcs , Basu:2005kt where energy–momentum fluctuations dominate the propagation and are bounded by analyzing cosmic-ray data. In our approach, energy violation is merely relevant for interactions of $\gamma$-rays with atmospheric nuclei as discussed here or with cosmic background fields. The foam impact on the latter reaction is up for debate, and, as already mentioned, is left for future study.
The main conclusion from this Letter, therefore, is that, in the string model of quantum foam considered here, subluminal modified photon dispersion relations (4) that may be compatible with a phenomenological hint of light-speed variation at a few $10^{17}$ GeV Shao:2009bv , Zhang:2014wpb , Amelino-Camelia:2016ohi , Xu:2016zxi , Xu:2016zsa , Xu:2018ien , Liu:2018qrg , Chen:2019avc , Zhu:2021pml , Zhu:2022usw , Song:2024and , Song:2025qej , Song:2025myx , Song:2025akr can evade drastic modifications to pair creation and bremsstrahlung under certain conditions. Thus, LV with such intensities may concordant in the context of such models with tests of subluminal photons in giant air showers. Otherwise the strong suppression effects induced in other LV proposals that are not observed experimentally in shower formation would allow Satunin:2023yvj to imply trans-Planckian limits for the relevant scale. A detailed estimate of more complicated effects of foam from dynamical aspect is anticipated to refine the above analysis, which might help better probing or constraining this type of QG models from string theory.
@@ Acknowledgments
A potential new line of evidence for the light-speed variation Shao:2009bv , Zhang:2014wpb , Amelino-Camelia:2016ohi , Xu:2016zxi , Xu:2016zsa , Xu:2018ien , Liu:2018qrg , Chen:2019avc , Zhu:2021pml , Zhu:2022usw , Song:2024and , Song:2025qej , Song:2025myx , Song:2025akr was recently reported Song:2025ksi , based on the detection of a 300 TeV photon-like air shower event by Carpet–3 coincident with GRB 221009A, offering further support for the theory here championed.